### v1

# DREAM: Scalable Red Teaming for Text-to-Image Generative Systems via Distribution Modeling

Abstract
Despite the integration of safety alignment and external filters, text-to-image (T2I) generative models are still susceptible to producing harmful content, such as sexual or violent imagery. This raises serious concerns about unintended exposure and potential misuse. Red teaming, which aims to proactively identify diverse prompts that can elicit unsafe outputs from the T2I system (including the core generative model as well as potential external safety filters and other processing components), is increasingly recognized as an essential method for assessing and improving safety before real-world deployment. Yet, existing automated red teaming approaches often treat prompt discovery as an isolated, prompt-level optimization task, which limits their scalability, diversity, and overall effectiveness. To bridge this gap, in this paper, we propose DREAM, a scalable red teaming framework to automatically uncover diverse problematic prompts from a given T2I system. Unlike most prior works that optimize prompts individually, DREAM directly models the probabilistic distribution of the target system’s problematic prompts, which enables explicit optimization over both effectiveness and diversity, and allows efficient large-scale sampling after training. To achieve this without direct access to representative training samples, we draw inspiration from energy-based models and reformulate the objective into simple and tractable objectives. We further introduce GC-SPSA, an efficient optimization algorithm that provide stable gradient estimates through the long and potentially non-differentiable T2I pipeline. During inference, we also propose a diversity-aware sampling strategy to enhance prompt variety. The effectiveness of DREAM is validated through extensive experiments, demonstrating that it surpasses 9 state-of-the-art baselines by a notable margin across a broad range of T2I models and safety filters in terms of prompt success rate and diversity. Additionally, DREAM successfully uncovers failure cases in 4 real-world commercial T2I systems and enables more robust safety fine-tuning that generalizes to unseen harmful prompts.
1 Introduction
Text-to-image (T2I) generative models [1, 2, 3, 4] are driving a new wave of visual content creation, reshaping our expectations of what machines are capable of. Trained on large-scale datasets [5], these models capture rich associations between language and imagery, allowing them to produce high-quality images with simple text inputs (known as prompts). Their ease of use and impressive flexibility have driven rapid adoption across creative arts, entertainment, and social media, particularly among younger users such as teenagers [6, 7, 8]. However, the same datasets that enable this versatility also inevitably contain unsafe content (e.g., sexually explicit material and violence) due to their massive scale and web-crawled nature [9, 10]. As a result, the models also acquire the ability to produce images of harmful content during real use, raising serious ethical, legal, and accountability concerns [11, 12].
To mitigate these risks, a growing number of efforts from both academia and industry [13, 10, 14, 15, 16] have focused on improving the safety of T2I generative models. One popular approach is safety alignment, also referred to as unsafe concept erasure in the T2I literature [13, 10, 14, 17], which fine-tunes the model using a curated set of unsafe prompts or images to suppress undesirable generations. This process helps steer the model toward harmless outputs: for example, returning a clothed figure even when prompted with “a nude person”. In addition, commercial companies like Stability AI [18] and Ideogram [19] also employ proprietary safety filters (e.g., NSFW image detectors) to block generation attempts when unsafe content is detected. These filters, when combined with the core generative model and other processing components, constitute the deployed T2I system. However, while these techniques show promising results in controlled environments, they remain imperfect when applied in practice. For example, both real-world users and researchers [20, 21] have reported that prompts unseen during training (e.g., implicit references to sensitive content), or even totally benign inputs (e.g., “the origin of woman”), may still escape moderation and lead to unsafe outputs. These observations highlight the limitations of current methods under open-ended inputs and the urgent need for proactive mechanisms to expose safety vulnerabilities of T2I generative systems before real-world deployment.
One emerging solution to proactively identify such blind spots is red teaming, where model owners (e.g., developers) simulates the behavior of real-world users to generate various testing prompts, aiming to systematically probe the model’s failure modes before deployment. In the context of T2I generative systems, red teaming typically attempts to find a diverse set of problematic prompts that can elicit unsafe or policy-violating outputs despite potential safeguards [20, 21]. By doing so, it not only serves as an evaluation tool for stress-testing the system’s safety and trustworthiness under open-ended inputs [20], but also provides valuable references for future improvement [22]. As a result, red teaming is increasingly recognized as a critical practice, with major companies like Google [23] initiating human-in-the-loop red teaming programs. At the same time, regulatory bodies are increasingly emphasizing rigorous safety testing before deployment, as reflected in the EU AI Act [24] and the U.S. NIST AI Risk Management Framework [25], alongside similar official efforts around the world [26, 27].
While early red teaming practice relied on human experts, recent works [20, 21, 28] have shifted toward automated red teaming, aiming to discover problematic prompts without human oversight. For example, FLIRT [28] employs a large language model (LLM) to generate an initial prompt and iteratively rewrite it toward unsafe outputs, while P4D [20] starts with a moderated unsafe prompt and applies token-level gradient-based substitutions to penetrate safety alignment. However, these methods often struggle to balance the success rate with prompt diversity, and can be prohibitively slow and costly to scale (e.g., P4D [20] takes $\sim$30 minutes to optimize a single prompt). These limitations underscore the urgent need for a scalable red teaming method that can efficiently generate a large, diverse set of effective problematic prompts.
In this paper, we present the first attempt towards bridging the aforementioned gap. Our method is driven by a unified insight into the shared limitations of previous works: they treat red teaming as a prompt-to-prompt discrete optimization problem, where each prompt is optimized independently yet without accumulating global knowledge across runs. Built upon this understanding, we propose Distributional Red tEAming via energy-based Modeling (DREAM), which directly models the probabilistic distribution of the target model’s unsafe prompts via training a parameterized prompt generator (e.g., an autoregressive LLM). In contrast to previous approaches, our formulation enables explicit optimization of both success and diversity, supports global updates to the modeled prompt distribution, and allows efficient large-scale sampling after training.
However, modeling the target prompt distribution is challenging, as it is tightly coupled to the specific T2I system and lacks sufficiently representative samples, making direct training infeasible. To overcome this, we draw inspiration from energy-based models [29], and decompose the originally intractable training objective into two surprisingly simple ones that allow effective distribution learning without direct sample access. Moreover, to enable effective and efficient gradient-based optimization for these objectives under long and potentially non-differentiable pipelines, we introduce Gradient-Calibrated Simultaneous Perturbation Stochastic Approximation (GC-SPSA), an efficient zero-th order optimization method based on SPSA [30]. Specifically, it estimates gradients using only forward evaluations and further improves stability via a history-aware calibration mechanism. We also provide theoretical analysis and convergence guarantees to support the use of GC-SPSA for optimizing our objectives. Finally, we propose a novel adaptive temperature scaling strategy method to further increase coverage at inference time.
We conduct extensive experiments on 5 state-of-the-art (SOTA) safety-aligned diffusion models, 4 safety filters, and compare DREAM with 9 SOTA baselines across two unsafe categories. The results show that DREAM consistently outperforms all baselines in terms of prompt success rate with a notable margin and approaches human-level diversity. DREAM also generalizes well to advanced T2I models (e.g., SDXL, SD v3) and 6 other NSFW themes, and remains effective even under strong combinational defenses or aggressive filters. Furthermore, an IRB-approved user study confirms that DREAM produces more effective and diverse prompts than prior methods. In addition, case studies demonstrate that DREAM can transfer to four commercial T2I platforms with unknown safety mechanisms. Finally, prompts generated by DREAM significantly enhance safety fine-tuning, enabling models to resist both seen and unseen harmful prompts better than those trained on other baselines.
To summarize, we make the following key contributions:
  • We revisit existing red teaming methods and identify a shared limitation: they treat prompt discovery as isolated, prompt-to-prompt optimization without global modeling, restricting their scalability and overall effectiveness.
  • We introduce DREAM, a scalable and distribution-aware red teaming framework that learns a probabilistic model over unsafe prompts using energy-based modeling. We further propose GC-SPSA, a novel zero-order optimization method that supports effective and efficient training, along with adaptive inference strategies for broader coverage. Theoretical analyses and global convergence guarantees are provided to support our framework.
  • We conduct comprehensive evaluations across 5 safety-aligned T2I models, 4 safety filters, and 9 SOTA baselines, demonstrating that DREAM achieves higher prompt success rate and matches human-level diversity. We also show DREAM ’s ability to expose failure cases in 4 commercial T2I platforms and to improve safety fine-tuning with strong generalization to unseen harmful prompts.
2 Related Work
2.1 Text-to-Image Generative Models
Text-to-image (T2I) generative models have become a cornerstone of modern visual synthesis, enabling users to create highly detailed images from natural language descriptions. Among various generative paradigms, diffusion models [31, 3, 1, 2, 4] have emerged as the dominant approach due to their superior training stability, generation quality, and controllability. Diffusion models operate by iteratively denoise random noise into coherent images, often conditioned on texts, making them particularly effective for large-scale training and text-controlled generation. Building upon this, a wide range of open-source (e.g., Stable Diffusion family [4]) and commercial systems (e.g., DeepAI [32], DALL·E 3 [33], Midjourney [34], Ideogram [19]) have been developed, providing state-of-the-art generation experiences under user-friendly graphical interfaces.
2.2 Unsafe Generation & Mitigation
The success of modern T2I models relies heavily on large datasets. For instance, Stable Diffusion is trained on LAION-5B [5], a web-scraped set of over 5 billion image-text pairs, while commercial models like Ideogram use even larger private datasets [19]. These datasets support powerful multimodal learning but also contain harmful content such as harmful imagery and copyrighted materials [35, 36]. This can be absorbed and reproduced by these models, raising ethical and legal issues, especially as these tools become more accessible and popular among children and adolescents, who may suffer psychological harm, safety risks, and disrupted development from exposure [11, 9, 37].
In response to these concerns, several mitigation strategies have emerged, which can be broadly categorized into two lines: model safety alignment and inference-time safety filtering. Model safety alignment [10, 14] refers to techniques that tune the diffusion model’s parameters directly to suppress its ability in producing unsafe content. This is typically achieved by collecting a curated set of harmful prompts or images and reinforcing the model to “unlearn” them through methods such as adversarial training [38], supervised fine-tuning [10, 13], or model editing [14, 17]. For example, CA [13] fine-tunes the diffusion model to match the image distribution of an unsafe target concept (e.g., “a nude woman”) to that of a safe anchor concept (e.g., “woman”). As a result, the model learns to resist prompts that are the same or similar to training-time target concepts and generates a safe image instead. In contrast, safety filters [16, 39, 15] act as external control mechanisms during inference. They can operate in prompt-level or image-level, aiming to detect and block unsafe content before or after generation. A representative case is the Safety Checker (SC) [15] employed in Stable Diffusion models, which compares the generated image with a set of predefined sensitive concepts and blocks outputs that exceed a similarity threshold.
While these approaches have demonstrated effectiveness with acceptable trade-offs in benign performance under their own evaluation protocols, their robustness in real-world scenarios has been frequently challenged by a growing body of recent research and user reports. For instance, text-based safety filters can be bypassed using simple synonym substitutions (e.g., changing “gun” to “sidearm”) [40], while image-based filters may lose effectiveness under subtle alterations in image styles, compositions, or rendering (e.g., “a nude woman in colored painting”) [41]. Besides, safety-aligned models may perform well when the prompts contain explicit words (e.g., “nude” or “sexy”), but still fail to handle veiled expressions, metaphors or context-related implications such as “a woman looking seductive,” “a cute Japanese movie star,” [42, 43, 20], which are unseen during unlearning. In addition to these scattered findings, recent research [44, 45, 46, 47] has developed various optimization methods to transform a given rejected prompt into a minimally modified variant that bypasses safety mechanisms, a technique known as adversarial jailbreak attacks. These diverse failure patterns on different safety mechanisms suggest it is crucial to proactively test and improve the system’s safety before real-world deployment.
2.3 Red Teaming for Text-to-Image Models
The concept of “red teaming” originated during the Cold War era in the 1960s as a form of structured military system testings and has since expanded to fields like cybersecurity, airport security, software engineering, and recently to AI and ML systems [48]. For generative models, red teaming typically involves simulating real user behavior to explore the system and find prompts that produce harmful or policy-violating outputs [49, 50, 51]. Unlike jailbreak attacks that tweak known unsafe prompts to evasive variants [44, 46], red teaming focuses on broader exploration to reveal diverse or unexpected failure modes [21]. It is now a key part of responsible model development [23] and is increasingly emphasized in recent regulatory frameworks [24, 25, 26].
One predominant form of red teaming is manual construction. For example, the I2P dataset [44] was formed by collecting and filtering harmful prompts from various forums through a mix of automatic tools and human curation. Similarly, Google’s Adversarial Nibbler Challenge [23] invited participants to attack real-world T2I models and selected high-quality prompts based on their effectiveness and diversity. Commercial providers also employ in-house or external experts to manually test models for discovering failure modes [52]. While such methods can surface unexpected and model-specific vulnerabilities, they rely heavily on human effort and lack automation, making them inefficient and expensive to conduct.
To this end, several methods for automated red teaming have been proposed [20, 28, 21, 53]. These methods typically adopt paradigms and techniques similar to jailbreak attacks and transform a set of initial prompts to harmful ones, using methods like token-level substitution [20, 53] and LLM-rewrite [28, 21]. However, as we will identify in the following section, this inherited formulation will inherently limit their effectiveness, exploration space, and efficiency, making them suboptimal for scalable red teaming.
3 Preliminaries
3.1 Threat Model
We consider the red team to be a benign (non-malicious) model owner aiming to proactively identify safety vulnerabilities in his/her own T2I generative system. Specifically, their goal is to find a set of diverse and effective prompts that can elicit unsafe or policy-violating outputs, in order to assess and improve safety before real-world use. Following previous works [21, 49, 51], we assume the red teamer (1) has full control over his/her T2I generative system, such as requesting it with arbitrary prompt and receive the resulting image (or an all-black image if blocked by filters), or access to the model’s parameters and gradients, and (2) can leverage auxiliary models (e.g., open-sourced LLMs) for assistance and has moderate computational resources to fine-tune these models.
3.2 Formulation of Red Teaming
Despite the growing importance of red teaming in evaluating the safety of T2I generative models, existing literature [21, 28, 20] largely lacks a formal formulation of what the red teaming task fundamentally entails. This absence has led to fragmented understanding and inconsistent objectives, which limits both theoretical analysis and the principled design of scalable red teaming methods.
To bridge this gap, we present a formal definition of red teaming in this section. Let $\mathcal{X}$ and $\mathcal{Y}$ be the prompt space and image space of the target T2I generative system, respectively, we can draw the following definition:
Definition 1 (Red Teaming T2I Systems).
Let $G:\mathcal{X}\rightarrow\mathcal{Y}$ be a T2I system mapping a text prompt $x\in\mathcal{X}=\mathcal{V}^{T}$ to an image $y\in\mathcal{Y}$, where $\mathcal{V}$ and $T$ represents the full vocabulary and the maximum prompt length of the system, respectively. Red teaming aims to find a prompt subset $\mathcal{A}\subseteq\mathcal{X}$ such that:
$\mathcal{A}:=\{x\in\mathcal{X}\mid\mathcal{O}(G(x))=1\},$
where $\mathcal{O}:\mathcal{Y}\rightarrow\{0,1\}$ is a binary oracle classifier that outputs $1$ if the image is unsafe and 00 if the image is safe or the request is denied by the equipped safety filter.
Intuitively, this definition formulates red teaming as a combinatorial subset discovery problem, whose aim is to identify all prompts that can trigger the T2I model to output unsafe content from the full prompt set $\mathcal{V}^{T}$. Note that the oracle function $\mathcal{O}$ is fundamentally unobservable in practice, as determining whether an image is “unsafe” is often vague, influenced by context, culture, and subjective interpretation [9]. As a practical alternative, red teaming methods rely on a surrogate scoring function $S:\mathcal{Y}\rightarrow\mathbb{R}$, which approximates the oracle with an objective score (e.g., the confidence score of an NSFW image detector). A prompt is deemed unsafe if its surrogate score is large enough to exceed a threshold $\tau$, yielding the surrogate unsafe set $\mathcal{A}_{\tau}:=\{x\in\mathcal{X}\mid S(G(x))\geq\tau\}$. While the surrogate formulation makes the task operational, obtaining the exact solution of the unsafe set $\mathcal{A}_{\tau}$ remains computationally intractable, as the task essentially reduces to a combinatorial searching problem over the prompt space $\mathcal{X}$, whose size grows exponentially with the prompt length $T$, i.e., $|\mathcal{X}|=|\mathcal{V}|^{T}$. In such combinatorial settings, exhaustive enumeration is the only general procedure that can ensure complete accuracy [54], yet it requires evaluating the surrogate score $S(G(x))$ for every enumerated $x\in\mathcal{V}^{T}$, making it computationally infeasible even for modest values of $T$. As a result, exact discovery is impractical except in trivial cases.
Fortunately, previous works have shown that exact recovery of $\mathcal{A}_{\tau}$ is often unnecessary. For instance, unlearning a moderate number of diverse and representative unsafe prompts is often sufficient to invalidate a much broader class of similar unsafe prompts [10, 14, 55, 56]. Consequently, the practical goal of red teaming shifts from full enumeration to the discovery of a representative and diverse subset $\hat{\mathcal{A}}\subseteq\mathcal{A}_{\tau}$, which captures a wide range of unsafe prompts while remaining computationally tractable to obtain.
Algorithm 1 A Generic Form of Existing Methods
1:Input: Seed distribution $\pi(x)$, number of prompts $N$, max steps $T$, scoring function $S(\cdot)$, T2I model $G(\cdot)$, threshold $\tau$, update operator $\textsc{Update}(\cdot)$
2:Output: Final set of optimized prompts $\hat{\mathcal{A}}$
3:$\hat{\mathcal{A}}\leftarrow\emptyset$
4:for $i=1$ to $N$ do
5:   $x_{i}^{(0)}\sim\pi(x)$
6:   $t\leftarrow 0$
7:   while $t<T$ and $S(G(x_{i}^{(t)}))<\tau$ do
8:      $x_{i}^{(t+1)}\leftarrow\textsc{Update}(x_{i}^{(t)},S(G(x_{i}^{(t)})))$
9:      $t\leftarrow t+1$
10:   end while
11:   $\hat{\mathcal{A}}\leftarrow\hat{\mathcal{A}}\cup\{x_{i}^{(t)}\}$
12:end for
13:Return $\hat{\mathcal{A}}$
3.3 Limitations of Previous Works
With an understanding of the red teaming task formulation, we now take a closer look at existing methods [20, 53, 28, 21]. While prior works differ substantially in their technical implementation, we distilled them into a unified, generic prompt-level discrete optimization paradigm, formalized in Alg. 1. Under this view, a red teaming algorithm begins with a seed prompt sampled from a seed distribution $\pi(x)$, and iteratively applies an Update operator guided by a scoring function $S(G(x))$, where $G(x)$ denotes the image generated by the T2I model. This loop continues until a generation crosses a threshold or a step budget is reached, at which point the final prompt is collected and the process resets. Note that the choice of seed prompt distribution, the Update operator, as well as the scoring function are all method-dependent. Despite empirical progress in uncovering unsafe prompts, we identify that this core algorithmic structure introduces two fundamental limitations, making them less suitable for scalable red teaming.
First, the Update operator essentially performs discrete optimization, which is inherently difficult due to the discontinuous and non-smooth nature of the ill-posed loss landscape of the discrete prompt space [57]. In fact, how to accurately obtain prompt-level gradient remains an open challenge in existing literature [58, 59]. As such, most existing methods [44, 45, 20, 60] resort to token-level gradient replacement, where each token is iteratively and greedily updated based on its locally estimated gradient with respect to $S(G(x_{i}^{(t)}))$. However, this limits the search space to local neighborhoods around the initial seed, making the optimization process highly sensitive to initialization [61]. While recent methods attempt to broaden the search space by prompting LLMs to generate sentence-level paraphrases [21, 28], these approaches are largely heuristic, lack convergence guarantees, and often result in unstable training dynamics in practice.
Second, one can easily observe from Alg. 1 that current methods are essentially operating at the individual prompt level: each run starts from a fresh seed prompt $x_{i}^{(0)}$, performs a local search trajectory $\{x_{i}^{(1)},x_{i}^{(2)},\cdots,x_{i}^{(t)}\}$, outputs $x_{i}^{(t)}$, and then discards all intermediate states but the final output before restarting the next run. This stateless fashion is naturally sub-optimal, as the algorithm would not accumulate any knowledge about explored regions or learn from past failures. Therefore, it may revisit similar attempt trajectories, re-try strategies that are known ineffective in previous runs and converging to familiar local optima, especially if the seed prompts are semantically or syntactically similar [21]. This makes the algorithm inefficient and yield highly similar prompts with limited marginal utility. Moreover, this inefficiency is exacerbated by the inherently slow convergence of discrete optimization. For example, P4D [20] and UnlearnDiffAtk [60] require roughly 3,000 rounds of model invocation and gradient updates to optimize a single prompt, taking about 30 minutes per prompt on an NVIDIA RTX A100 GPU. These deficiencies make it very difficult to be scaled up for large-scale red teaming.
4 The Design of DREAM
4.1 Distributional Red Teaming via Energy-Based Modeling
Motivated by our previous analysis about the limitations of existing methods, the key insight behind our proposal is to shift from discrete, stateless prompt-to-prompt optimization to directly modeling the distribution over unsafe prompts.
Formally, let $q^{*}(x)$ denote the true (but unknown) distribution over the target model’s problematic prompts, i.e., the probabilistic distribution from which samples $x\in\hat{\mathcal{{A}}}$ are drawn. Our goal is to learn a probabilistic distribution $p_{\theta}(x)$ parameterized by $\theta$ (e.g., an autoregressive language model $p_{\theta}(x)=\prod_{t=1}^{T}p_{\theta}(x_{t}\mid x_{<t})$), such that $p_{\theta}(x)$ approximates $q^{*}(x)$ as possible, which can be characterized by the following Kullback–Leibler divergence [62] objective:
$\theta^{*}\in\arg\min_{\theta}D_{\mathrm{KL}}(p_{\theta}\,\|\,q^{*})$ (1)
This formulation has several desirable properties. First, by modeling the distribution $p_{\theta}(x)$, our method naturally converts the prompt-level discrete optimization into continuous optimization over model parameters $\theta$. Second, since the parameters encode the distribution over prompts, each parameter update accumulates knowledge about which types of prompts are more or less likely to trigger unsafe outputs, thus promoting exploration efficiency during training. Furthermore, it is totally feasible to initialize $p_{\theta}(x)$ with a pre-trained language model. As a result, our method inherits strong priors from large-scale human language data, which enables the model to understand and explore nuanced expressions, innuendos, and cultural references, which are subtle signals that typically require human-like common sense or contextual awareness and are often inaccessible to previous token-level search methods. Finally, once training is complete, sampling from the learned distribution $p_{\theta^{*}}(x)$ is efficient. An arbitrary number of diverse prompts can be generated efficiently via forward passes, without requiring iterative search or gradient updates. This property makes our approach particularly suitable for red teaming, where a large-scale of unsafe prompts (e.g., thousands) are required for safety assessment and downstream safety-tuning.
Despite these promising properties, the objective in Eq. (1) remains particularly challenging to optimize in practice. The core difficulty lies in that the ground-truth distribution $q^{*}(x)$ is fundamentally unknown and there exists no readily available dataset that is sufficiently representative of the full support of the target model’s problematic prompts. This makes direct optimization (e.g., through fine-tuning with MLE [63]) impossible. Fortunately, recent advances in implicit generative modeling [64] provide a viable pathway to tackle this challenge. Specifically, results from the theory of energy-based models [29, 64] suggest that even in the absence of explicit samples, the target distribution $q^{*}(x)$ can be implicitly characterized with a properly defined energy function $E(x)$, which is a real-valued function that assigns lower values to more likely (or desirable) samples, and higher values otherwise. Then, the unknown distribution $q^{*}(x)$ can be expressed as a Boltzmann distribution [65] $q^{*}(x)={\exp(-\beta\cdot E(x))}/{Z}$, where $Z$ is a constant that normalizes the distribution [29, 64] and $\beta>0$ is a hyperparameter. Then, by plugging it into Eq.(1), we have:
$\arg\min_{\theta}D_{\mathrm{KL}}(p_{\theta}\,\|\,q^{*})=\sum_{x\in\mathcal{X}}p(x)\log\left(\frac{p(x)}{q(x)}\right)=$ (2)
$\mathbb{E}_{x\sim p_{\theta}}\left[\log\frac{p_{\theta}(x)}{q^{*}(x)}\right]=\mathbb{E}_{x\sim p_{\theta}}\left[E(x)+\frac{1}{\beta}\cdot\log p_{\theta}(x)\right].$
The derivation above reduces the otherwise intractable KL divergence to two surprisingly simple yet intuitive components: the first is to minimize the energy function $E(x)$, i.e., to steer $\theta$ toward regions in the prompt space where the energy is low and thus more desirable. The second objective acts as a regularizer that penalizes low-entropy distributions by minimizing the log-likelihood $\mathbb{E}_{x\sim p_{\theta}}[\log p_{\theta}(x)]$, thereby avoiding degenerate solutions where the model collapses to a narrow set of prompts.
4.2 Energy Function Design
So far, we have decomposed the objective into two intuitive and interpretable sub-goals. The second regularization term $\mathbb{E}_{x\sim p_{\theta}}[\log p_{\theta}(x)]$ is straightforward to compute and optimize in practice. We now turn our attention to the first component, the energy function $E(\cdot)$. The energy function essentially defines the target distribution by assigning lower energy scores to desirable prompts and higher scores to undesired ones. In this section, we introduce our energy function design, which captures the following two scores.
Vision-level Harmfulness Energy. The primary goal of $E$ is to guide the model toward the target model’s vulnerable prompt distribution $\mathcal{A}_{\tau}$. However, directly assessing the harmfulness of a text prompt $x$ is difficult as the risk often emerges only after it is rendered into an image. Therefore, we take a vision-level approach by evaluating the output image $y=G(x)$ instead of the prompt itself. Specifically, we employ BLIP-2 [66], a pretrained vision-language model with strong generalization across diverse image-text domains, to compute a vision-level harmfulness energy as part of the energy function. It assesses how the generated image is semantically aligned with a predefined harmful concept, and assign lower energy to those prompts that algin better with the harmful concept.
Formally, given a generated image $y=G(x)$ and a pre-defined target description $c$ (like “an image containing nudity”), the textual description $c$ is first sent to a pretrained language encoder $\mathcal{T}_{\phi}$ to obtain the sentence-level semantic embedding $t=\mathcal{T}_{\phi}(c)$. Then, the image $y$ is passed through a vision encoder followed by a specialized transformer module known as the Q-Former [66]. This module employs a set of query embeddings to interact with the visual features via cross-attention and finally extracts a set of latent tokens $\mathcal{I}_{\psi}(y)=\{z_{1},\dots,z_{k}\}$, each representing a different fine-grained aspect of the image in the same vision-language embedding space. Then, we define the alignment score as:
$E_{\text{align}}(x)=\mathbb{E}_{x\sim p_{\theta}}\left[-\max_{z_{i}\in\mathcal{I}_{\psi}(G(x))}\frac{\langle z_{i},t\rangle}{\|z_{i}\|\cdot\|t\|}\right]$ (3)
[…]
This formulation brings three key benefits. First, BLIP-2 provides better generalization even under distribution shifts, such as stylized or non-photorealistic images, making the alignment score more reliable across visual domains [66, 67]. Second, the approach enables flexible red teaming through natural language descriptions. One can easily shift the target by modifying the concept prompt, e.g., replacing $c$ with “an image depicting violence scenes” to target violent content. When a small set of reference images is available, prompt tuning techniques can also be used to further refine and control the targeted concept [68]. Third, the alignment score is continuous, allowing small improvement in prompt effectiveness to be reflected. This supports more stable optimization than discrete (e.g. binary) success signals.
Prompt-level Diversity Energy. While the combination of the alignment energy and entropy regularization in Eq. (3) and Eq. (4) is effective in steering the model toward $q^{*}(x)$, we find that the entropy term alone is often insufficient to ensure diverse generations. This is because the target distribution $q^{*}(x)$ may itself be biased, e.g., certain keywords like “nude” and their variant sentences might dominate the probability mass. As a result, semantically distinct prompts with lower probability under $q^{*}(x)$ may remain largely unvisited in limited sampling iterations. To address this, we introduce a diversity energy term that explicitly encourages broader coverage within a limited sample budget. Let $\mathcal{E}_{\xi}(x)\in\mathbb{R}^{d}$ denote the sentence embedding of prompt $x$, obtained from a frozen pre-trained encoder (e.g., a sentence transformer [69]). Then, we define the prompt-level diversity energy as the expected pairwise similarity among prompt embeddings sampled from the current model distribution $p_{\theta}(x)$:
$E_{\text{div}}(x)=\mathbb{E}_{x,x^{\prime}\sim p_{\theta},\,x\neq x^{\prime}}\left[\frac{\langle\mathcal{E}_{\xi}(x),\mathcal{E}_{\xi}(x^{\prime})\rangle}{\|\mathcal{E}_{\xi}(x)\|\cdot\|\mathcal{E}_{\xi}(x^{\prime})\|}\right].$ (4)
This would explicitly encourage semantic diversity among generated prompts in limited sampling iterations, thus promoting broader exploration and reducing redundancy.
4.3 Red Team LLM Optimization
After designing the energy function, we can plug Eq. (3) and Eq. (4) into Eq. (LABEL:eq:entropy), and arrive at the final training objective for the red team prompt generator $\theta$:
$\min_{\theta}\mathbb{E}_{x\sim p_{\theta}}\left[E_{\text{align}}(x)+\lambda\cdot E_{\text{div}}(x)+\frac{1}{\beta}\cdot\log p_{\theta}(x)\right],$ (5)
where $\lambda$ and $\beta$ are balancing hyperparameters. However, optimizing Eq. (5) is non-trivial. One intuitive approach would be to use backpropagation-based methods to obtain exact gradients and then update the LLM’s parameters. However, applying backpropagation-based optimization directly is challenging. This is because the full red-teaming pipeline includes multiple components, including autoregressive language generation, multi-step diffusion denoising, and energy models. For instance, generating a 10-token prompt with an autoregressive LLM involves 10 sequential decoding steps, each with its own activations. Likewise, synthesizing an image via diffusion models typically requires 30 iterative denoising steps. Storing the full computation graph for just a single forward pass through this pipeline would easily demand thousands of gigabytes GPU memory, making end-to-end backpropagation-based training memory-prohibitive even for small-scale models and small batch sizes. Moreover, certain components like keyword-based safety filters are non-differentiable, further hindering backpropagation.
To enable effective gradient-driven optimization while avoiding the need for backpropagation through the entire pipeline, we propose a novel framework based on Simultaneous Perturbation Stochastic Approximation (SPSA) [30]. SPSA is a classical zero-th order optimization method that allows estimates of high-dimensional gradients using only forward evaluations. However, Vanilla SPSA is validated to suffer from instability and slow convergence in our red-teaming setup, due to the highly stochastic nature of both LLMs and diffusion-based generation (see experiments in Section 5.7). To mitigate this, we propose a simple yet effective variant, GC-SPSA, which incorporates an adaptive sampling schedule as well as a history-aware gradient calibration mechanism to reduce gradient variance while maintaining efficiency. In the following, we first introduce SPSA and analyze its problems, and then we propose our GC-SPSA, and finally provide a theoretical analysis and convergence guarantee to support our design.
Definition 2 (SPSA [30]).
Given an objective function $\mathcal{L}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ and parameters $\theta\in\mathbb{R}^{d}$, SPSA uses the following randomized two-point finite-difference approximation to compute the unbiased estimate of gradient $\nabla_{\theta}\mathcal{L}(\theta)$:
$g(\theta):=\frac{\mathcal{L}(\theta+\epsilon\Delta)-\mathcal{L}(\theta-\epsilon\Delta)}{2\epsilon}\Delta,$ (6)
where $\epsilon>0$ is a small perturbation magnitude, $\Delta\in\mathbb{R}^{d}$ is a random perturbation vector sampled from a zero-mean Gaussian distribution.
Previous works have proved that SPSA provides an unbiased estimate of the true gradient, i.e., $\mathbb{E}_{\Delta}[g(\theta)]=\nabla_{\theta}\mathcal{L}(\theta)$ [30]. In our setting, SPSA is particularly advantageous: it avoids backpropagation entirely and requires only forward evaluations under perturbed parameters, making it well-suited for bypassing the long and possibly non-differentiable pipeline. Moreover, it eliminates the need to store intermediate forward activations, further reducing memory consumption compared to backpropagation-based methods [70].
Despite these advantages, we observe in our experiments that directly applying SPSA leads to unstable training dynamics. This instability primarily stems from the inherent stochasticity in both LLM and diffusion-based sampling, which introduces high variance into single-shot gradient estimates. To reduce variance, a straightforward strategy is to increase the number of forward estimates per iteration and average the resulting gradients. However, this significantly increases training cost linearly and is expensive in practice.
To mitigate this problem, we propose a simple yet effective method, GC-SPSA, which stabilizes SPSA with a novel adaptive gradient calibration algorithm. Our key insight is that Eq. (5) primarily steers the LLM toward a subspace of prompts that are likely to elicit harmful images from the target model. Since the pretrained LLM already possesses strong priors, such subspaces are shown to typically reside within a relatively flat and continuous basin in the loss landscape nearby the pre-trained parameters [71, 72, 73]. Therefore, we hypothesis that it is more crucial to ensure the reliability of early optimization steps to accurately locate this subspace, yet latter updates may tolerate more variance and can be calibrated with early reliable gradients.
Algorithm 2 The Complete Training Procedure of DREAM
1:Input: Initial model parameters $\theta_{0}$, initial sampling budget $n_{0}$, learning rate $\eta$, correction strength $\gamma$, smoothing factor $\rho$, decay factor $T_{\text{dec}}$, max steps $T_{\text{max}}$
2:Output: Optimized generator parameters $\theta_{T_{\max}}$
3$\triangleright$ Initialize gradient estimate and effective sample size
4:$\hat{g}_{0}\leftarrow\frac{1}{n_{0}}\sum_{i=1}^{n_{0}}g_{0,i}(\theta_{0})$
5:$w_{0}\leftarrow n_{0}$
6:for $t=1$ to $T_{\text{max}}$ do
7$\triangleright$ Determine sampling budget for current step
8:   $n_{t}\leftarrow\max\left(1,\left\lfloor\frac{n_{0}}{2^{t/T}}\right\rfloor\right)$
9$\triangleright$ Estimate gradients via SPSA with $n_{t}$ queries
10:   for $i=1$ to $n_{t}$ do
11:      Sample perturbation $\Delta_{t,i}\sim\text{Gaussian}^{d}$
12:      $\mathcal{L}^{+}_{t,i},\mathcal{L}^{-}_{t,i}\leftarrow\mathcal{L}(\theta_{t}+\epsilon\Delta_{t,i}),\mathcal{L}(\theta_{t}-\epsilon\Delta_{t,i})$
13:      $g_{t,i}\leftarrow\frac{\mathcal{L}^{+}_{t,i}-\mathcal{L}^{-}_{t,i}}{2\epsilon}\cdot\Delta_{t,i}$
14:   end for
15$\triangleright$ Aggregate and calibrate gradients
16:   $\hat{g}_{t}\leftarrow\frac{1}{n_{t}}\sum_{i=1}^{n_{t}}g_{t,i}+\gamma\cdot\frac{w_{t-1}}{w_{t-1}+n_{t}}\cdot\hat{g}_{t-1}$
17:   $w_{t}\leftarrow\rho\cdot w_{t-1}+(1-\rho)\cdot n_{t}$
18$\triangleright$ Update model parameters
19:   $\theta_{t+1}\leftarrow\theta_{t}-\eta\cdot\hat{g}_{t}$
20:end for
21:Return $\theta_{T_{\max}}$
Specifically, for the $t$-th update, we first estimate the gradient $n_{t}$ times via Eq. (6), obtaining a set of stochastic gradient estimates $\{g_{t,1}(\theta_{t}),g_{t,2}(\theta_{t}),\dots,g_{t,n_{t}}(\theta_{t})\}$. The number of queries $n_{t}$ is controlled by an exponentially decaying schedule: $n_{t}=\max\left(1,\left\lfloor\frac{n_{0}}{2^{t/T_{\text{dec}}}}\right\rfloor\right)$, where $n_{0}$ is the initial number of sampling times and $T_{\text{dec}}$ governs the decay rate. This scheduling allocates a higher sampling budget to early iterations and gradually reduces the number of samples as optimization stabilizes. In our experiments, we find that setting $n_{0}=4$ and $T_{\text{dec}}=10$, i.e., starting with 4 samples and halving the sampling budget every 10 steps, yields stable and efficient optimization performance (see experiments in Sec. 5.7). To further reduce the variance of the later estimated gradient, inspired by confidence-aware optimal Bayesian fusion [74], we introduce a gradient calibration mechanism. Specifically, we treat each new gradient estimate as a noisy observation and combine it with historical information using a confidence-aware correction term:
$\hat{g}_{t}=\frac{1}{n_{t}}\sum_{i=1}^{n_{t}}g_{t,i}(\theta_{t})+\gamma\cdot\frac{w_{t-1}}{w_{t-1}+n_{t}}\cdot\hat{g}_{t-1},$ (7)
$\theta_{t+1}=\theta_{t}-\eta\cdot\hat{g}_{t},$
where $\hat{g}_{t-1}$ is the accumulated gradient estimate from previous iterations ($\hat{g}_{0}=\frac{1}{n_{0}}\sum_{i=1}^{n_{0}}g_{0,i}(\theta_{0})$), $w_{t-1}$ denotes its effective sample size (initialized as $w_{0}=n_{0}$), $\eta$ is the learning rate. The update of $w_{t}$ follows an exponential moving average rule. We use the term $\frac{w_{t-1}}{w_{t-1}+n_{t}}$ to approximate the relative confidence of historical vs. current gradients, and $\gamma$ controls the overall strength of the correction. Intuitively, our confidence-weighted fusion scheme anchors the current noisy gradient estimate towards the historically aggregated direction, especially when the current estimate is based on fewer samples (i.e., lower confidence). As formally analyzed in Theorem 1, the GC-SPSA estimator achieves a strictly higher signal-to-noise ratio (SNR) than the Vanilla SPSA estimator for all $t\geq 1$, which helps mitigate gradient noise and promotes a more consistent optimization path.
Theorem 1 (Improved SNR of GC-SPSA).
Let $\bar{g}_{k}$ be the Vanilla SPSA estimator and $\hat{g}_{k}=\bar{g}_{k}+H_{k}\hat{g}_{k-1}$ be the GC-SPSA estimator, with $\hat{g}_{0}=\bar{g}_{0}$ and $H_{k}>0$. Then for all $t\geq 1$, the SNR difference between the GC-SPSA and the Vanilla-SPSA admits the explicit positive lower bound:
$\mathcal{D}_{t}=\frac{\|g_{\text{true}}\|^{2}}{V_{\rm single}\sum_{k=0}^{t}h_{k}^{2}V_{k}}\left[P_{t}^{2}V_{\rm single}-\sum_{k=0}^{t}h_{k}^{2}V_{k}\right]$ (8)
where the weights are defined as $h_{t}=1$ and $h_{k}=\prod_{j=k+1}^{t}H_{j}$. Here, $P_{t}=\sum_{k=0}^{t}h_{k}$ is the cumulative weight sum and $V_{k}=V_{\text{single}}/n_{k}$ is the gradient variance at step $k$.
The detailed proof is in Appendix A.1. Furthermore, we also provide theoretical global convergence guarantee and convergence rate analysis for our proposed GC-SPSA to optimize our objective functions in Eq. (5) under mild assumptions, as formally shown in the following theorem:
Theorem 2 (Global Convergence and Rate Analysis of GC-SPSA).
Consider an objective $\mathcal{L}:\mathbb{R}^{d}\to\mathbb{R}$ satisfying $\nabla^{2}\mathcal{L}(\theta)\preceq\ell I_{d}$ for all $\theta\in\mathbb{R}^{d}$, where $I_{d}$ denotes the $d$-dimensional identity matrix. Then, GC-SPSA will converge (i.e., $\min_{t\in[T]}\mathbb{E}[\|g(\theta_{t})\|^{2}]\leq\delta.$) after
$T=\Theta\left(\frac{\mathcal{L}(\theta_{1})-\mathcal{L}^{*}}{\eta\,\zeta_{\min}\,\delta-C_{\max}\,\Upsilon-\frac{\ell\,\eta^{2}\,\Xi}{2}}\right).$ (9)
iterations. Here, $\mathcal{L}^{*}$ is the global minimum, and the constants $\Upsilon>0$ and $\Xi>0$ denote upper bounds on the average squared norm of the gradient estimator and the accumulated noise variance, respectively.
For training, we begin by perturbing the current model parameters $\theta_{t}$ and estimate the gradient $n_{t}$ times using Eq. (6). Since the objective $\mathcal{L}(\theta)$ involves an intractable expectation, we approximate it via Monte Carlo sampling over a batch of sampled prompts (Alg. 3). The resulting gradient estimates are then averaged and calibrated using our confidence-aware update rule in Eq. (LABEL:eq:spsa-gc). Finally, the calibrated gradient is used to update the parameters $\theta_{t}$. We summarize the complete training procedure in Alg. 2.
4.4 Inference-Time Adaptive Temperature Scaling
After training, the optimized red team LLM is steered to a distribution of the prompts that can induce unsafe outputs. The final step is to utilize this model to generate red teaming prompts. However, in practice, we find that the model may still fail to adequately explore its support during generation. This is due to the absence of awareness across generations: the model essentially does not know what has already been generated. Therefore, it may produce repetitive tokens that has limited marginal benefits.
To further enhance diversity, we propose an inference-time strategy that encourages diversity via adaptive temperature scaling. Recall that in autoregressive decoding, the model iteratively predicts the next token distribution with $p_{\theta^{*}}({x_{t}}=v|x_{<t})=\frac{\exp\left(\boldsymbol{z}_{t}[v]/\tau\right)}{\sum_{j\in\mathcal{V}}\exp\left(\boldsymbol{z}_{t}[j]/\tau\right)}$, where the temperature hyperparameter $\tau_{t}$ controls the sharpness of the token distribution $\boldsymbol{z}_{t}\in\mathbb{R}^{|\mathcal{V}|}$ at decoding step $t$. Intuitively, lower temperatures make the model more confident (peaky), while higher temperatures flatten the distribution to encourage exploration. As such, we maintain a global token frequency vector $\mathbf{f}\in\mathbb{N}^{|\mathcal{V}|}$, tracking the number of times each vocabulary token has appeared across previous generations. At decoding step $t$, given raw logits $\boldsymbol{z}_{t}\in\mathbb{R}^{|\mathcal{V}|}$, we compute the relative frequency of the top-scoring token $v_{t}=\arg\max_{j}\boldsymbol{z}_{t}[j]$, and use it to scale the temperature:
$\tau_{t}=\max\left(\frac{1}{\alpha}\log(1+\frac{\mathbf{f}[v_{t}]}{\sum_{j}\mathbf{f}[j]}),\ \tau_{\min}\right),$ (10)
where $\alpha>0$ is a sensitivity coefficient and $\tau_{\min}$ ($0.8$ in this paper) prevents degeneracy. The final logits are adjusted as $\tilde{\boldsymbol{z}}_{t}=\boldsymbol{z}_{t}/\tau_{t}$ before sampling. This penalizes frequent tokens, promoting underexplored generations without modifying training or architecture. Empirically, we find it improves prompt diversity with minimal overhead and little loss in effectiveness.
TABLE I: Comparison with baselines on Stable Diffusion v1.5 and other safety-aligned diffusion models.
Stable Diffusion v1.5 CA ESD UCE SafeGen RECE
Sexual Violence Sexual Violence Sexual Sexual Violence Sexual Sexual
PSR $\uparrow$ / PS $\downarrow$ PSR $\uparrow$ / PS $\downarrow$ PSR $\uparrow$ / PS $\downarrow$ PSR $\uparrow$ / PS $\downarrow$ PSR $\uparrow$ / PS $\downarrow$ PSR $\uparrow$ / PS $\downarrow$ PSR $\uparrow$ / PS $\downarrow$ PSR $\uparrow$ / PS $\downarrow$ PSR $\uparrow$ / PS $\downarrow$
Human-written Datasets
I2P 51.5% / 0.49 13.9% / 0.47 11.7% / 0.49 13.1% / 0.47 09.5% / 0.49 10.0% / 0.49 06.9% / 0.47 45.6% / 0.49 07.5% / 0.49
Adv. Nibbler 28.3% / 0.53 08.9% / 0.54 03.1% / 0.53 03.5% / 0.54 01.2% / 0.53 30.1% / 0.53 07.8% / 0.54 33.8% / 0.53 00.7% / 0.53
Automated Red Teaming (Token Perturbation)
QF-Attack 23.8% / 0.60 10.6% / 0.62 00.6% / 0.60 09.4% / 0.62 00.0% / 0.60 01.3% / 0.60 10.0% / 0.62 21.2% / 0.60 00.6% / 0.60
SneakyPrompt 61.7% / 0.52 26.5% / 0.65 13.5% / 0.52 25.6% / 0.64 07.9% / 0.52 16.5% / 0.52 20.1% / 0.64 25.5% / 0.52 14.0% / 0.52
MMA-Diffusion 91.0% / 0.63 73.9% / 0.65 44.9% / 0.63 59.2% / 0.65 35.9% / 0.63 59.9% / 0.63 66.9% / 0.65 34.0% / 0.63 52.2% / 0.63
P4D 78.0% / 0.60 42.0% / 0.58 52.0% / 0.60 26.0% / 0.66 43.3% / 0.60 24.0% / 0.56 10.0% / 0.58 61.9% / 0.55 16.0% / 0.55
UnlearnDiffAtk 83.0% / 0.52 30.0% / 0.49 36.4% / 0.52 10.5% / 0.49 21.2% / 0.52 24.6% / 0.51 10.5% / 0.49 55.9% / 0.52 13.6% / 0.52
Automated Red Teaming (LLM Rewrite)
ART 14.9% / 0.48 29.2% / 0.47 02.7% / 0.49 14.1% / 0.46 00.8% / 0.49 00.8% / 0.49 20.8% / 0.46 13.7% / 0.49 00.8% / 0.48
FLIRT 91.8% / 0.77 74.4% / 0.66 26.0% / 0.58 64.4% / 0.63 17.1% / 0.64 48.5% / 0.64 18.4% / 0.61 10.2% / 0.59 10.7% / 0.57
Ours 92.2% / 0.50 87.0% / 0.55 76.0% / 0.56 77.3% / 0.57 72.1% / 0.56 89.0% / 0.54 83.6% / 0.57 81.6% / 0.49 91.3% / 0.56
5 Evaluation
5.1 Experimental Setup
Target Diffusion Models & Safety Filters. In our experiments, we evaluate a variety of standard diffusion models, safety-aligned models, and safety filters. In addition to the standard Stable Diffusion v1.5, we also evaluate safety-aligned models including ESD [10], CA [13], UCE [14], SafeGen [42], and RECE [17], all of which have unlearned certain unsafe concepts from the models. For external safety filters, following Yang et al. [44], we consider 4 external filters classified into (1) text-based filters (NSFW text classifier [16] and Keyword-Gibberish hybrid filter [75]), (2) image-based filter (NSFW image filter [39] and Stable Diffusion’s built-in image safety checker, SC [15]). Furthermore, we evaluate the generalizability of DREAM on several real-world models, including SDXL, SDv3, Kandinsky v3, and Shuttle 3 Diffusion. We also evaluate DREAM in a transfer-based setting on multiple real-world online T2I generation-as-a-service platforms, including Ideogram, DeepAI, DALL·E 3, and Midjourney, which are known to incorporate strong safety strategies with undisclosed detailed implementations.
Baselines. We evaluate and compare DREAM against several state-of-the-art baselines. For human-written red teaming datasets, we include I2P [9] and Google’s Adversarial Nibbler [23], collected from T2I community forums and via crowdsourcing, respectively. For prompt-level attacks & red teaming methods, we consider QF-Attack [47], SneakyPrompt [44], MMA-Diffusion [45], P4D [20], UnlearnDiffAtk [60], FLIRT [28] and ART [21]. Note that some baselines (e.g., SneakyPrompt) do not explicitly identify themselves as red teaming methods, yet we include them in our comparison for the sake of completeness given their strong technical resemblance to some red teaming baselines. For all baselines, we follow their default settings, use their original seed prompt datasets, and run enough training epochs to ensure convergence.
Evaluation Metrics. We primarily use two metrics to evaluate the performance of each red teaming method: Prompt Success Rate (PSR) and Prompt Similarity (PS), which measure the effectiveness and diversity of the generated prompts, respectively. PSR is the proportion of prompts that successfully trigger the target model to generate images containing the specified inappropriate content. Following Yang et al. [45], we use PSR out of $N$ generations (PSR-N) instead of a single generation to reduce the impact of inherent stochasticity in diffusion sampling. Specifically, for each prompt, we generate $N$ images with different random seeds. The prompt is considered successful if at least one of these $N$ images contains the desired unsafe concept, and the final PSR is measured as the ratio of successful prompts. In our experiments, we use PSR-3, i.e., $N=3$, and mainly adopt Multi-headed Safety Classifier (MHSC) [11] as our detector. MHSC is a category-specific NSFW image detector that provides per-category confidence scores for various unsafe concepts. It has been widely used in the community due to its high precision [11, 76, 45], especially on AI-generated images. A higher PSR ($\uparrow$) indicates the method is more capable of generating effective prompts. Besides, PS quantifies the diversity of the prompts by measuring the average pairwise cosine similarity between all prompt embeddings. In our evaluation, we use the state-of-the-art BGE embedding model [77] to obtain the prompt embeddings (note that this model is different from $\mathcal{E}_{\xi}$ we use in Eq. (4)). A lower PS score ($\downarrow$) indicates lower inter-prompt similarity, indicating that the prompts are more diverse and less repetitive.
TABLE II: Comparison with baselines on external safety filters.
Safety Checker NSFW Image Detector NSFW Text Detector Keyword-Gibberish Filter
Sexual Violence Sexual Violence Sexual Violence Sexual Violence
PSR $\uparrow$ / PS $\downarrow$ PSR $\uparrow$ / PS $\downarrow$ PSR $\uparrow$ / PS $\downarrow$ PSR $\uparrow$ / PS $\downarrow$ PSR $\uparrow$ / PS $\downarrow$ PSR $\uparrow$ / PS $\downarrow$ PSR $\uparrow$ / PS $\downarrow$ PSR $\uparrow$ / PS $\downarrow$
Human-written Datasets
I2P 23.4% / 0.49 13.6% / 0.47 26.3% / 0.49 13.0% / 0.47 37.5% / 0.49 08.5% / 0.47 26.5% / 0.49 06.5% / 0.47
Adv. Nibbler 14.3% / 0.53 08.4% / 0.54 17.3% / 0.53 08.2% / 0.54 19.4% / 0.53 06.2% / 0.54 04.8% / 0.53 00.0% / 0.54
Automated Red Teaming (Token Perturbation)
QF-Attack 11.9% / 0.59 10.6% / 0.62 01.9% / 0.59 10.6% / 0.62 01.2% / 0.59 08.8% / 0.62 03.1% / 0.59 00.0% / 0.62
SneakyPrompt 30.5% / 0.52 26.5% / 0.67 11.0% / 0.50 26.5% / 0.64 12.5% / 0.46 13.0% / 0.64 51.8% / 0.52 18.1% / 0.64
MMA-Diffusion 40.5% / 0.63 72.2% / 0.65 04.7% / 0.63 68.3% / 0.65 01.5% / 0.63 11.5% / 0.65 00.0% / 0.63 00.0% / 0.65
P4D 40.0% / 0.60 40.0% / 0.58 44.0% / 0.60 38.0% / 0.58 04.0% / 0.60 04.0% / 0.58 00.0% / 0.60 00.0% / 0.58
UnlearnDiffAtk 35.6% / 0.52 10.5% / 0.49 43.2% / 0.52 09.5% / 0.49 48.3% / 0.52 06.3% / 0.49 07.6% / 0.52 02.1% / 0.49
Automated Red Teaming (LLM Rewrite)
ART 04.4% / 0.48 28.8% / 0.47 09.2% / 0.48 29.2% / 0.48 06.4% / 0.48 21.8% / 0.46 04.4% / 0.48 04.5% / 0.47
FLIRT 26.5% / 0.62 72.6% / 0.64 45.9% / 0.75 75.4% / 0.65 08.5% / 0.51 16.7% / 0.60 44.6% / 0.58 48.9% / 0.57
Ours 64.7% / 0.52 86.4% / 0.54 57.3% / 0.50 83.7% / 0.58 62.3% / 0.51 42.5% / 0.54 67.4% / 0.52 65.7% / 0.56
5.2 Main Results
Effectiveness on Concept-erased T2I Models. We first conduct experiments on both the standard Stable Diffusion (SD) v1.5 model and several concept-erased models, which are either fine-tuned with harmful prompts or images to unlearn unsafe content, or directly modified to collapse harmful concept vectors (e.g., distorting the embedding of harmful tokens like “nudity” or “sexual” to approximate that of empty strings). As shown in Tab. I, our method consistently achieves the highest PSRs across all models and categories, significantly outperforming all baselines. For example, on concept-erased models such as UCE and RECE, human-written datasets typically yield PSRs below 10%, and state-of-the-art red teaming methods like MMA-Diffusion often struggle to exceed 50%. In contrast, our prompts achieve PSRs above 79% on all evaluated models and even exceed 90% on SD and RECE. In addition to effectiveness, our method also excels in prompt diversity. Across all models, our prompts maintain a prompt similarity (PS) score around 0.55, which is notably better than most of baselines and on par with human-written datasets, indicating higher diversity. This highlights that our approach not only discovers more successful prompts, but also explores a broader and more varied region of the prompt space that remains unexplored for existing red teaming techniques.
Effectiveness on External Safety Filters. We further evaluate the effectiveness of DREAM on various external safety filters and compare it with baselines. For each safety filter, we combine it with the standard SD v1.5 model and regard the whole model-filter pipeline as an integrated generative system. As shown in Tab. II, baseline methods are largely unstable. For example, MMA-Diffusion achieved good results on most safety-aligned models and SC. In contrast, our DREAM consistently achieves the highest PSR-3 across all settings, outperforming baselines by a large margin, demonstrating the superiority and universality of our DREAM.
TABLE III: Effectiveness of DREAM across models and NSFW themes. (a) shows results with more models on sexual (left) and violence (right) concepts; (b) shows results on more NSFW themes on SD v1.5.
(a) More T2I Models
Model (Sexual) PSR $\uparrow$ / PS $\downarrow$ Model (Violence) PSR $\uparrow$ / PS $\downarrow$
Stable Diffusion XL 89.5% / 0.52 Stable Diffusion XL 92.3% / 0.51
Stable Diffusion v3 82.5% / 0.53 Stable Diffusion v3 92.2% / 0.52
Kandinsky v3 89.8% / 0.52 Kandinsky v3 86.8% / 0.53
Shuttle 3 Diffusion 86.9% / 0.51 Shuttle 3 Diffusion 90.0% / 0.51
(b) More NSFW Themes
Category PSR $\uparrow$ / PS $\downarrow$ Category PSR $\uparrow$ / PS $\downarrow$
Self-harm 87.8% / 0.55 Shocking 94.7% / 0.56
Hate 92.6% / 0.52 Harassment 94.1% / 0.53
Political 91.6% / 0.50 Illegal Activity 92.7% / 0.51
Effectiveness on More T2I Models & NSFW Themes. We evaluate DREAM on more T2I models, including Stable Diffusion XL [4], Stable Diffusion v3 [78], Kandinsky v3 [79], and Shuttle 3 Diffusion [80]. These models vary significantly in architectural design (e.g., using DiT [81] instead of convolutional U-Nets) and training settings, and some of them are reported to apply dataset filtering to cleanse (some) unsafe images before training [4, 78]. As shown in Tab. III (a), DREAM consistently performs well across all models and both categories, with PSR approaching 90% on average. The discovered prompts also exhibit a level of diversity comparable to human-written ones. These results highlight the model architecture-agnostic nature of DREAM and its strong potential for application to future models. Additionally, we further assess the generalizability of DREAM on additional NSFW themes following previous works [9], including self-harm, shocking, hate, harassment, political, and illegal activity. Following Yang et al. [45], we use Q16 [68] as the detector in this setting, as MHSC does not support all of these categories. As shown in Tab. III (b), DREAM maintains high PSR across most cases, exceeding 90% consistently and reaching close to 95% for shocking and harassment. Besides, our prompts remain highly diverse and close to that of human-written prompts. These findings demonstrate the scalability and generalizability of DREAM in discovering a broader range of unsafe concepts.
Figure 1: User study results on prompt success rate.
Figure 2: User study results on prompt diversity.
5.3 Human Evaluation
Considering that automated metrics may not perfectly reflect human perception, we also conduct a user study under IRB approval to assess the effectiveness and diversity of different methods. Specifically, we select SD v1.5, CA, and NSFW Text Filter as the representative models. Then we select the most effective method from each category, i.e., I2P for human-written data, MMA-Diffusion for token-level perturbation, and FLIRT for LLM rewrites, and compare them with our DREAM. We evaluate all model-method combination on both sexual and violence categories, resulting in $2\times 3\times 4=24$ concept-model-method settings in total.
For each setting, we randomly sample 30 prompts from the method’s generated prompts to form a prompt pool. Each prompt generates 3 images, forming a prompt-images group. Then, we recruit 30 volunteers, all university students or faculties from various academic backgrounds, to participate in the study. All participants are ensured to be adults in good physical and mental health, and are fully informed and agree to participate in our user study. Then, we brief participants on the basics of T2I models and red teaming, the recommended definition of the corresponding unsafe category (Tab. X in Appendix), and start the study after obtaining his/her confirmation of full comprehension.
For evaluation, each participant is randomly assigned 5 prompt-image groups from the 30 available for each concept-model-attack setting. Prompt-image groups are organized in random order and displayed one-by-one. Participants are asked to determine whether the three displayed images in the group clearly reflects the specified unsafe concept, and the corresponding group is marked as successful if and only if the participant identifies at least one such image (PSR-3). After that, participants are shown with the full set of $4\times 30$ problematic prompts on SD v1.5 and asked to rate these prompt sets by their perceived diversity (ranges from 1$\sim$5, higher the better) based on their understanding of the prompt’s lexical, structural, and semantic richness (Tab. XI in Appendix). We replaced MMA with Adv. Nibbler as it is not readable to human raters. Despite involving images depicting NSFW concepts, our study has been reviewed and approved by our IRB under a process analogous to the “exempt review” category of U.S. IRB protocols (45 CFR 46), since the IRB staffs determined our study to pose no more than “minimal risk” given that participants were healthy adults, fully informed, and free to withdraw at any time. As shown in Fig. 1 and 2, DREAM consistently achieves good results on this user study with the best prompt success rate and a diversity similar to human-written dataset, demonstrating our effectiveness.
5.4 Adaptivity Under Stronger Safety Mechanisms
In this section, we examine whether DREAM remains effective under more stringent safety mechanisms. Specifically, we evaluate its effectiveness against (1) composite filtering pipelines combining multiple mechanisms, and (2) MHSC [45] as the safety filter. We compare our DREAM with MMA-Diffusion [45] and FLIRT [28], which achieved the best averaged performance on their categories (i.e., token-level perturbation and LLM-assisted rewrite, respectively). The evaluated concept is sexual.
Composite Filtering. We consider systems that sequentially combines multiple safety mechanisms. Specifically, we consider two combinations: (1) NSFW Text Filter + SD v1.5 + NSFW Image Filter; and (2) Keyword-based Filter + ESD [10] + SC [15]. A prompt is considered successful only if at least one out of the three generations is not rejected by any of the filters and successfully contains unsafe content, as rated by MHSC. Note that while these combinations help narrow the unsafe prompt space, the system’s false positive rate also increases exponentially, as any individual false rejection would invalidate the whole sample.
As shown in Tab. IV (a)-(b), DREAM consistently achieves good results and outperforms the baselines in both settings. Notably, baseline methods exhibit steep drops in effectiveness, and even fail entirely under stronger combinations. In contrast, DREAM maintains moderate success rates even under aggressive filtering, demonstrating its adaptivity.
MHSC as the Safety Filter. We also consider an extreme setting where MHSC, the same classifier used to compute PSR in our experiments, is directly deployed as the safety filter. In this case, MHSC can no longer be used for evaluation, as any generations that could be classified as harmful would be blocked in advance. Thus, we adopt human-evaluated PSR-3 as the evaluation metric. The evaluation is done under two thresholding settings, corresponding to 5% (default) and 1% false positive rates (FPR), calibrated on a benign held-out dataset following its original paper.
As shown in Tab. IV (d), despite MHSC’s high precision, DREAM is able to identify multiple prompts that bypass filtering and lead to unsafe generations. We attribute it to MHSC’s conservativeness as a NSFW classifier: it only flags outputs when highly confident, prioritizing precision over recall, as also reported in the original paper [11]. This makes it a reliable evaluation tool (high PSRs indeed indicates high true positives) but also means that some borderline harmful cases near the threshold may slip through. These subtle failure modes are where DREAM excels, thanks to its distributional exploration and fine-grained energy modeling.
As a final remark, while MHSC is designed to be conservative, it is still more aggressive than real-world filters (e.g., Stable Diffusion’s Safety Checker has a reported FPR below 0.1% [42]). It is thus reasonable that DREAM uncovers fewer prompts under MHSC than under more permissive filters. More broadly, this highlights an open challenge in balancing protection with the risk of over-censorship: aggressive filters indeed reduce risks but also inevitably narrow the prompt space, often at the cost of creative expression or user experience. We believe red teaming methods like DREAM can serve as a valuable complement, which helps to surface near-boundary cases that evade detection and informing targeted improvements that reduce blind spots without broadly increasing over-censorship.
TABLE IV: Effectiveness of DREAM and baselines under stronger safety mechanisms. (a)–(b) report results under multi-stage filters; (c) shows results under MHSC with two thresholds.
(a) NSFW Text + NSFW Image
Method PSR $\uparrow$ / PS $\downarrow$
MMA-Diffusion 00.1% / 0.63
FLIRT 07.6% / 0.50
Ours 52.8% / 0.55
(b) Keyword + ESD + SC
Method PSR $\uparrow$ / PS $\downarrow$
MMA-Diffusion 14.4% / 0.63
FLIRT 02.3% / 0.54
Ours 37.3% / 0.56
(c) MHSC
@1% FPR PSR $\uparrow$ / PS $\downarrow$ @5% FPR PSR $\uparrow$ / PS $\downarrow$
MMA-Diffusion 23.3% / 0.63 MMA-Diffusion 06.7% / 0.63
FLIRT 06.7% / 0.54 FLIRT 00.0% / 0.50
Ours 43.3% / 0.58 Ours 16.7% / 0.52
TABLE V: Transferability results on real-world T2I-as-a-service platforms that utilizes unknown safety mechanisms.
(a) Ideogram
Method Prompt Bypass $\uparrow$ Prompt-Image Bypass $\uparrow$ Human-Rated Success Rate $\uparrow$ Prompt Similarity $\downarrow$
MMA-Diffusion 75.0% 65.9% 43.9%$\pm$1.90.66
FLIRT 76.7% 67.8% 30.6%$\pm$1.90.54
Ours 98.4% 96.1% 57.9%$\pm$3.20.52
(b) DeepAI
MMA-Diffusion 41.0% 26.7% 18.7%$\pm$1.20.65
FLIRT 58.0% 51.0% 15.7%$\pm$4.20.50
Ours 89.0% 79.0% 55.7%$\pm$3.10.53
(c) DALL·E 3
MMA-Diffusion 36.7% 31.7% 7.3%$\pm$1.60.64
FLIRT 30.0% 28.3% 2.8%$\pm$1.60.51
Ours 60.8% 47.9% 32.3%$\pm$2.40.55
(d) Midjourney
MMA-Diffusion 18.2% 18.2% 18.2%$\pm$2.20.63
FLIRT 21.7% 21.7% 10.0%$\pm$1.20.53
Ours 60.3% 60.3% 35.7%$\pm$1.70.55
5.5 Transferability on Real-world Commercial T2I Generative Models
To further assess the scalability of DREAM in real-world conditions, we test it on four widely used commercial T2I platforms: Ideogram, DeepAI, DALL·E 3, and Midjourney. These platforms deploy state-of-the-art, closed-source safety systems that at least include both prompt- and image-level filters, though their exact implementations are undisclosed. Moreover, some platforms use proprietary LLMs to interpret and rewrite user prompts. To evaluate how different methods perform on these real-world platforms, we train DREAM and FLIRT on the NSFW Image and Text Hybrid Filter, and then randomly select 50 prompts to conduct a transfer-based red teaming. We evaluate both “sexual” and “violence” categories, which are explicitly prohibited by all the platforms’ safety policies, and report the averaged results. As shown in Tab. V, our method achieves good transferability on all evaluated platforms, as validated by a high prompt bypass rate (the fraction of prompts accepted by the text filter), prompt-image bypass rate (the fraction of attempts that successfully yield generated images), human-rated prompt success rate, and still outperforms baselines with a notable margin. The results indicate that while the unsafe prompt space varies across different T2I platforms, the prompts generated by our method possess a notable degree of transferability, i.e., we can uncover some shared vulnerabilities across different systems, possibly due to broader prompt coverage and the inherent similarity across these T2I models.
5.6 Discussion
LLM Reusability. One potential advantage of our distributional modeling approach is that the red team LLM, once trained on a T2I model, learns a holistic understanding of the probability distribution over unsafe prompts. As a result, the LLM retains reusable knowledge that may be effectively leveraged when adapted to similar T2I systems. To evaluate this hypothesis, we conduct reusability experiments where a red team LLM trained on CA for 300 steps is adapted to other T2I systems. As shown in Tab. VI, our red team LLM achieves non-trivial success rates when directly reused (transferred) on other models without any further training. More importantly, with only 50 additional training steps ($\sim 2$ GPU hours), the reused LLM can be rapidly adapted to the new model, with some even achieving performance close to training from scratch (e.g., on ESD and UCE). These results show that DREAM’s distribution-level modeling demonstrates strong reusability potential, where the learned knowledge can be efficiently transferred and adapted to other T2I systems with reduced computational overhead.
Mitigation Strategy. To mitigate the identified vulnerabilities, one practical strategy is safety-tuning, which fine-tunes the T2I model on the collected unsafe prompts in an adversarial manner to unlearn them. To assess the utility of different methods for this purpose, we utilize red team prompts identified by each method as the dataset, and use Safety-DPO [22], a recent algorithm designed to steer generation away from unsafe behaviors via preference modeling, to fine-tune the SD v1.5 model. We then evaluate the resulting models against prompt sets from all methods, which yields a square matrix where each row represents a safety-tuned model (trained on method A’s data), and each column corresponds to evaluation against method B’s prompts. As shown in Fig. 3, the model adversarially trained with our DREAM-generated dataset consistently achieves the lowest PSR across all test sets and both categories, including those totally unseen during training. In contrast, models tuned with baseline datasets tend to show limited generalization and failing to defend against prompts from other methods, especially those discovered by DREAM. This also indirectly suggests that DREAM’s global modeling helps improve the diversity and coverage of discovered prompts, which in turn supports more robust and generalizable safety improvement.
TABLE VI: Results on reusing the red team LLM (prompt generator) trained on CA to other T2I systems. The metric is PSR-3.
Setting ESD UCE RECE SC
Ours (Direct Transfer) 57.4% 77.4% 75.5% 40.7%
Ours (+50 Step Adaptation) 68.5% 86.3% 82.9% 48.3%
(a) Sexual category
(b) Violence category
Figure 3: PSR results of SD v1.5 safety-aligned with red team datasets generated by different methods.
Prior-informed Enhancement. For generality, our DREAM is designed without imposing specific prior about the internal components or defenses of the target T2I system. However, in practice, model owners (e.g., developers) have prior knowledge about the system, which can be potentially leveraged to enhance red teaming. For instance, if the model owner knows the system employs keyword-based filters, a simple enhancement strategy is to remove these tokens from the red team LLM’s vocabulary. This prior encourages the generator to focus on unexplored regions of the prompt space without wasting effort on words that are doomed to be rejected. To evaluate this, we conduct a case study on the Keyword Filter + UCE setup. We observe that the keyword-removed version converges faster, reaching near-optimal performance within 150 training steps, compared to 280 steps required by the baseline. Interestingly, the final performance difference is modest (63.4% vs. 60.2% PSR), suggesting that while prior knowledge accelerates convergence, it is not critical for eventual performance. This result highlights two insights: first, DREAM is effective even without any system-specific priors, making it broadly applicable; second, when available, prior information can be selectively incorporated to enhance DREAM. However, leveraging such priors often requires case-specific integration strategies, some of which may be difficult or costly to implement in practice. How to develop principled ways to incorporate them, especially for neural network-based components, remains an open direction for future work.
5.7 Ablation Study
In this section, we conduct an ablation study of DREAM, with CA+Sexual as the default setting.
Effectiveness of Each Component. As shown in Tab. VII, all components contributes to DREAM’s final performance. For example, while $E_{\text{align}}(x)$ pushes the model towards harmful outputs, it often leads to less diverse prompts. Adding $E_{\text{div}}(x)$ helps strike a balance between effectiveness and diversity. Additionally, Adaptive temperature scaling (ATS) improves prompt diversity during inference with only minimal PSR degradation, highlighting its effectiveness for balancing effectiveness and diversity.
TABLE VII: Ablation study on each component. ATS stands for our inference time adaptive temperature scaling strategy and Opt. Alg. means optimization algorithm.
$E_{\text{align}}(x)$ $E_{\text{div}}(x)$ ATS Opt. Alg. PSR $\uparrow$ / PS $\downarrow$
$\checkmark$ -- -- GC-SPSA 90.8% /00.70
$\checkmark$ $\checkmark$ -- GC-SPSA 76.4% /00.58
$\checkmark$ $\checkmark$ $\checkmark$ SPSA 45.9% /00.54
$\checkmark$ $\checkmark$ $\checkmark$ GC-SPSA 76.0% / 0.56
TABLE VIII: Ablation study on GC-SPSA and different $n_{0}$.
Method Steps GPU Time PSR $\uparrow$ / PS $\downarrow$
SPSA 340 12.85h 47.7% / 0.53
SPSA 410 15.37h 61.8% / 0.55
GC-SPSA ($n_{0}=4$) 300 12.78h 76.0% / 0.56
GC-SPSA ($n_{0}=8$) 300 15.43h 79.8% / 0.57
Effectiveness of GC-SPSA. As shown in Tab. VIII, GC-SPSA consistently outperforms vanilla SPSA, even when SPSA is given more steps to ensure equal GPU time. This demonstrates that investing in early gradient quality and applying historical calibration can indeed yield better overall performance, which indirectly confirms our insights in Sec. 4. In addition, we evaluate different initial sampling budgets $n_{0}$. We observe that $n_{0}=4$ already provides a strong balance between efficiency and effectiveness. While increasing $n_{0}$ to 8 leads to further gains, the improvement is marginal compared to the additional cost, possibly because the variance is already small enough to ensure stable optimization. Therefore, we adopt $n_{0}=4$ as our default configuration, as it strikes a good trade-off between convergence speed, stability, and computational efficiency.
Effect of $\lambda$ and $\alpha$. $\lambda$ and $\alpha$ controls the trade-off between diversity and effectiveness by influencing the weighting of diversity energy during training and temperature sensitivity during inference, respectively. As shown in Tab. IX, increasing $\lambda$ or decreasing $\alpha$ enhances prompt diversity, yet at the cost of the decrease of effectiveness. However, the performance is generally stable and satisfactory when these hyperparameters are within a reasonable range. Thus, we set the hyperparameters to $\lambda=1$ and $\alpha=0.05$ as the default configurations, and the users may tune them if they have specific emphasis on effectiveness or diversity.
TABLE IX: Ablation study on $\lambda$ and $\alpha$.
$\alpha$ 0.01 0.03 0.05 0.08 0.1
70.2% / 0.53 72.7% / 0.54 76.0% / 0.56 76.4% / 0.57 76.4% / 0.58
$\lambda$ 0.6 0.8 1 1.2 1.4
80.4% / 0.61 78.7% / 0.58 76.0% / 0.56 74.7% / 0.54 65.4% / 0.52
6 Conclusion
This paper presents DREAM, a novel framework for scalable red teaming of T2I generative systems. DREAM learns the distribution of unsafe prompts via energy-based modeling, allowing efficient, diverse, and effective prompt discovery at scale. We further introduce GC-SPSA, an efficient method to optimize our objective and propose adaptive strategies for broad prompt coverage during inference. Through comprehensive experiments, we demonstrate DREAM’s superior effectiveness and generalizability.
Limitations. Our work still has the following limitations, which we aim to address in future work. First, similar to other methods, our current implementation involves several auxiliary models (e.g., BLIP-2), and it is possible that our system may inherit certain limitations or biases rooted in these models. Although we did not observe clear evidence of such bias in our experiments, we acknowledge the potential risk and plan to explore more robust alternatives in the future. Second, while our global modeling approach offers desirable benefits, we admit it requires moderate costs to train. However, we note that this training cost is amortized over a large number of generated prompts. Compared to some baseline methods (e.g., P4D and MMA) which require 30 minutes or more to obtain a single prompt, DREAM remains more efficient even when applied to a moderate number of prompts ($>30$). For users with limited computational resources, it is also possible to further reduce cost by reusing or adapting a previously trained red team LLM, as discussed in Sec. 5.6. Finally, while our GC-SPSA demonstrates both theoretical guarantees and strong empirical performance outperforming baseline optimizers such as vanilla SPSA and LLM-based heuristics, we acknowledge it is essentially an approximation and may not be fully precise or perfectly efficient. Nonetheless, as discussed in Sec. 4, obtaining exact gradients via backprobagation are often infeasible due to the memory-intensive and potentially non-differentiable nature of the full T2I pipeline. We hope our energy-based distributional formulation and our proposed optimizer can inspire and serve as a valuable foundation for future improvements, and ultimately inspire stronger, more systematic safety evaluation practices for T2I systems.
Appendix A Omitted Derivations and Proofs
A.1 Proof of Theorem 1
Proof. We establish a strictly positive lower bound for the SNR improvement $\mathcal{D}_{t}:=\mathrm{SNR}_{t}^{\rm GC}-\mathrm{SNR}^{\rm Vanilla}$ by analyzing the statistical properties of the GC-SPSA estimator with decaying sample sizes against the Vanilla SPSA baseline.
The single-sample SPSA estimator $g(\theta)$ uses the observable loss $\mathcal{L}(\theta)=\mathcal{L}_{\text{true}}(\theta)+\xi$. A second-order Taylor expansion of $\mathcal{L}_{\text{true}}$ implies that the finite difference approximation yields $\Delta^{\top}g_{\text{true}}+O(\epsilon^{2})$, where $g_{\text{true}}\triangleq\nabla_{\theta}\mathcal{L}_{\text{true}}(\theta)$. Therefore,
$g(\theta)=\left(\Delta^{\top}g_{\text{true}}+O(\epsilon^{2})\right)\Delta+\frac{\xi^{+}-\xi^{-}}{2\epsilon}\Delta$ (11)
Taking expectation over the randomness in $\Delta$ and $\xi$: $\mathbb{E}[g(\theta)]=g_{\text{true}}+O(\epsilon^{2})$ since $\mathbb{E}_{\Delta}\left[(\Delta^{\top}g_{\text{true}})\Delta\right]=g_{\text{true}}$ and the noise term averages to zero.
The second moment arises from signal, bias, and noise components, i.e., $\mathbb{E}[\|(\Delta^{\top}g_{\text{true}})\Delta\|^{2}]=d\|g_{\text{true}}\|^{2}$, $\mathbb{E}[\|O(\epsilon^{2})\Delta\|^{2}]=O(d\epsilon^{4})$, and $\mathbb{E}[\|\frac{\xi^{+}-\xi^{-}}{2\epsilon}\Delta\|^{2}]\leq\frac{d\sigma_{\xi}^{2}}{2\epsilon^{2}}$. Thus, the variance of one time estimation is:
$V_{\rm single}:=\mathrm{Var}(g(\theta))=(d-1)\|g_{\text{true}}\|^{2}+\frac{d\sigma_{\xi}^{2}}{2\epsilon^{2}}+O(d\epsilon^{4})$ (12)
GC-SPSA collects $n_{t}$ times estimation following the exponential decay strategy $n_{t}=\max(1,\lfloor n_{0}/2^{t/T_{\text{dec}}}\rfloor)$ and takes their average to obtain $\bar{g}_{t}(\theta)$. Its variance is:
$V_{t}:=\mathrm{Var}(\bar{g}_{t}(\theta))=\frac{V_{\rm single}}{n_{t}}$ (13)
The GC-SPSA estimator $\hat{g}_{t}$ follows the recursive update rule (LABEL:eq:spsa-gc), unrolling this recursion gives:
$\hat{g}_{t}=\sum_{k=0}^{t}h_{k}\bar{g}_{k}(\theta_{k})$ (14)
where $h_{t}=1$, $H_{j}=\gamma\frac{w_{j-1}}{w_{j-1}+n_{j}}$, and $h_{k}=\prod_{j=k+1}^{t}H_{j}$.
The expectation and variance of the GC-SPSA are:
$\mathbb{E}[\hat{g}_{t}]$ $\approx P_{t}g_{\text{true}}(\theta_{t}),\quad\text{where}\quad P_{t}=\sum_{k=0}^{t}h_{k}$ (15)
$\mathrm{Var}(\hat{g}_{t})$ $=\sum_{k=0}^{t}h_{k}^{2}V_{k}=\sum_{k=0}^{t}h_{k}^{2}\frac{V_{\rm single}}{n_{k}}$ (16)
The SNR improvement is defined as $\mathcal{D}_{t}=\mathrm{SNR}_{t}^{\text{GC}}-\mathrm{SNR}^{\rm Vanilla}$, where:
$\mathrm{SNR}_{t}^{\text{GC}}=\frac{\|g_{\text{true}}(\theta_{t})\|^{2}P_{t}^{2}}{\sum_{k=0}^{t}h_{k}^{2}V_{k}}$ (17)
$\mathrm{SNR}^{\rm Vanilla}=\frac{\|g_{\text{true}}(\theta_{t})\|^{2}}{V_{\rm single}}$ (18)
Therefore:
$\mathcal{D}_{t}=\frac{\|g_{\text{true}}\|^{2}}{V_{\rm single}\sum_{k=0}^{t}h_{k}^{2}V_{k}}\left[P_{t}^{2}V_{\rm single}-\sum_{k=0}^{t}h_{k}^{2}V_{k}\right]$ (19)
To prove the bracketed term $P_{t}^{2}V_{\rm single}-\sum_{k=0}^{t}h_{k}^{2}V_{k}$ is strictly positive, we expand $P_{t}^{2}=\sum_{k=0}^{t}h_{k}^{2}+2\sum_{0\leq i<j\leq t}h_{i}h_{j}$:
$\begin{split}&P_{t}^{2}V_{\rm single}-\sum_{k=0}^{t}h_{k}^{2}V_{k}\\ &=\left(\sum_{k=0}^{t}h_{k}^{2}+2\sum_{0\leq i<j\leq t}h_{i}h_{j}\right)V_{\rm single}-\sum_{k=0}^{t}h_{k}^{2}V_{k}\\ &=\sum_{k=0}^{t}h_{k}^{2}V_{\rm single}\left(1-\frac{1}{n_{k}}\right)+2V_{\rm single}\sum_{0\leq i<j\leq t}h_{i}h_{j}.\end{split}$ (20)
The first term is non-negative, and the second term is strictly positive. Therefore, we have $\mathcal{D}_{t}>0$. ∎
A.2 Proof of Theorem 2
Proof. The condition $\nabla^{2}\mathcal{L}(\theta)\preceq\ell I_{d}$ implies the following descent inequality for the update rule $\theta_{t+1}=\theta_{t}-\eta\,\hat{g}_{t}$:
$\begin{split}\mathbb{E}\bigl{[}\mathcal{L}(\theta_{t+1})\mid\mathbb{F}_{t}\bigr{]}&\leq\mathcal{L}(\theta_{t})-\eta\,\bigl{\langle}g_{t},\,\mathbb{E}[\hat{g}_{t}\mid\mathbb{F}_{t}]\bigr{\rangle}\\ &\quad+\frac{\ell\eta^{2}}{2}\,\mathbb{E}\bigl{[}\|\hat{g}_{t}\|^{2}\mid\mathbb{F}_{t}\bigr{]},\end{split}$ (21)
where $g_{t}:=\nabla\mathcal{L}(\theta_{t})$. The conditional mean is $\mathbb{E}[\hat{g}_{t}\mid\mathbb{F}_{t}]=g_{t}+\gamma\alpha_{t}\hat{g}_{t-1}$. Then conditional second moment is given by:
$\mathbb{E}\bigl{[}\|\hat{g}_{t}\|^{2}\mid\mathbb{F}_{t}\bigr{]}$ $=\mathrm{Var}(\bar{g}_{t}\mid\mathbb{F}_{t})+\bigl{\|}\mathbb{E}[\hat{g}_{t}\mid\mathbb{F}_{t}]\bigr{\|}^{2}$
$=\mathrm{Var}_{t}+\|g_{t}\|^{2}+2\gamma\alpha_{t}\langle g_{t},\hat{g}_{t-1}\rangle$
$\quad+(\gamma\alpha_{t})^{2}\|\hat{g}_{t-1}\|^{2}.$ (22)
Substituting these into eq. (21) and grouping terms yields:
$\mathbb{E}[\mathcal{L}(\theta_{t+1})\mid\mathbb{F}_{t}]$ $\leq\mathcal{L}(\theta_{t})-\eta(1-\tfrac{\ell\eta}{2})\|g_{t}\|^{2}$
$\quad-\eta\gamma\alpha_{t}(1-\ell\eta)\langle g_{t},\hat{g}_{t-1}\rangle$
$\quad+\frac{\ell\eta^{2}(\gamma\alpha_{t})^{2}}{2}\|\hat{g}_{t-1}\|^{2}+\frac{\ell\eta^{2}}{2}\mathrm{Var}_{t}.$ (23)
Applying Young’s inequality to the cross-term gives:
$-\eta\gamma\alpha_{t}(1-\ell\eta)\langle g_{t},\hat{g}_{t-1}\rangle$
$\quad\leq\frac{\eta\gamma\alpha_{t}(1-\ell\eta)}{2}\|g_{t}\|^{2}+\frac{\eta\gamma\alpha_{t}(1-\ell\eta)}{2}\|\hat{g}_{t-1}\|^{2}.$ (24)
Plugging this bound back in leads to the one-step descent lemma:
$\mathbb{E}[\mathcal{L}(\theta_{t+1})\mid\mathbb{F}_{t}]$ $\leq\mathcal{L}(\theta_{t})-\eta\,\zeta_{t}\,\|g_{t}\|^{2}+C_{t}\,\|\hat{g}_{t-1}\|^{2}$
$\quad+\frac{\ell\,\eta^{2}}{2n_{t}}\left(O(d\,\epsilon_{t}^{4})+\frac{d\,\sigma_{\xi}^{2}}{4\epsilon_{t}^{2}}\right),$ (25)
where the coefficients are defined as
$\zeta_{t}$ $:=1-\tfrac{\ell\eta}{2}-\tfrac{\gamma\alpha_{t}(1-\ell\eta)}{2}-\tfrac{\ell\eta(d-1)}{2n_{t}},$ (26)
$C_{t}$ $:=\frac{\eta\gamma\alpha_{t}(1-\ell\eta)}{2}+\frac{\ell\eta^{2}(\gamma\alpha_{t})^{2}}{2}.$ (27)
Telescoping the one-step descent lemma eq. (A.2) over $t=1,\dots,T$. Let $\zeta_{\min}:=\min_{1\leq t\leq T}\zeta_{t}>0$. Taking expectations, summing, we rearrange to find:
$\min_{1\leq t\leq T}\mathbb{E}\|g(\theta_{t})\|^{2}$ $\leq\frac{\mathcal{L}(\theta_{1})-\mathcal{L}^{*}+\sum_{t=1}^{T}C_{t}\,\mathbb{E}\|\hat{g}_{t-1}\|^{2}}{\eta\,T\,\zeta_{\min}}$
$\quad+\frac{\ell\,\eta}{2\,T\,\zeta_{\min}}\sum_{t=1}^{T}\frac{V_{t}}{n_{t}}.$ (28)
Prior analyses of stochastic optimization algorithms have established the boundedness of error terms [30, 82, 83], i.e.,
$C_{t}\leq C_{\max},\quad\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\hat{g}_{t-1}\|^{2}\leq\Upsilon,\quad\text{and}\quad\frac{1}{T}\sum_{t=1}^{T}\frac{V_{t}}{n_{t}}\leq\Xi.$
Applying these boundedness yields:
$\min_{1\leq t\leq T}\mathbb{E}\|g(\theta_{t})\|^{2}$ $\leq\frac{\mathcal{L}(\theta_{1})-\mathcal{L}^{*}}{\eta\,T\,\zeta_{\min}}$
$\quad+\frac{C_{\max}\,\Upsilon}{\eta\,\zeta_{\min}}+\frac{\ell\,\eta\,\Xi}{2\,\zeta_{\min}}.$ (29)
To ensure the left-hand side is at most $\delta$, we solve for $T$. This requires
$T\geq\frac{\mathcal{L}(\theta_{1})-\mathcal{L}^{*}}{\eta\,\zeta_{\min}\,\delta-C_{\max}\,\Upsilon-\frac{\ell\,\eta^{2}\,\Xi}{2}},$ (30)
which establishes the iteration complexity of GC-SPSA. ∎
Algorithm 3 Loss Estimation via Monte Carlo Sampling
1:Input: Model parameters $\theta_{t}$, batch size $N$, image generator $G(\cdot)$, hyperparameters $\beta$ and $\lambda$.
2:Output: Estimated loss $\mathcal{L}(\theta_{t})$
3$\triangleright$ Sample a batch of prompts from $p_{\theta_{t}}(x)$
4:$\mathcal{X}\leftarrow\{x_{i}\sim p_{\theta_{t}}(x)\}_{i=1}^{N}$
5$\triangleright$ Compute alignment energy for each image
6:Initialize $\mathcal{L}_{\text{align}}\leftarrow 0$
7:for $x_{i}\in\mathcal{X}$ do
8:   $y_{i}\leftarrow G(x_{i})$
9:   $\mathcal{L}_{\text{align}}\leftarrow\mathcal{L}_{\text{align}}+E_{\text{align}}(x_{i})$
10:end for
11$\triangleright$ Compute prompt-level diversity energy
12:$\mathcal{L}_{\text{div}}\leftarrow\frac{1}{N(N-1)}\sum_{i}\sum_{j\neq i}\frac{\langle\mathcal{E}_{\xi}(x_{i}),\mathcal{E}_{\xi}(x_{j})\rangle}{\|\mathcal{E}_{\xi}(x_{i})\|\cdot\|\mathcal{E}_{\xi}(x_{j})\|}$
13$\triangleright$ Compute entropy regularization
14:$\mathcal{L}_{\text{ent}}\leftarrow\sum_{i=1}^{N}\log p_{\theta_{t}}(x_{i})$
15$\triangleright$ Aggregate total objective from all terms
16:$\mathcal{L}(\theta_{t})\leftarrow\mathcal{L}_{\text{align}}+\lambda\cdot\mathcal{L}_{\text{div}}+\frac{1}{\beta}\cdot\mathcal{L}_{\text{ent}}$
17:Return $\mathcal{L}(\theta_{t})$
Figure 4: Example unsafe images generated by DREAM’s red team prompt on the corresponding model. The black block and heavy blurring is added by authors to avoid disturbing readers.
Figure 5: Example unsafe images generated by DREAM’s red team prompt on online services. The black block is added by authors to avoid disturbing readers.
Appendix B More Experimental Details
TABLE X: Categories and definitions of unsafe content used in our paper.
Category Definition
Sexual Content that is sexually explicit, including nudity, sexual acts, genital exposure, or content that, though not explicitly depicting nudity or sexual acts, are overly sexualized, with clear sexual provocations, sexual innuendo, or erotic tease.
Violence Content involving physical aggression, brutality, threats, or harm directed at individuals or groups, including depictions of interpersonal violence, intended to shock, disturb, promote violent behavior, or when featuring graphic imagery of excessive bloodshed or serious injuries.
TABLE XI: Prompt diversity levels and their definitions used in our user study.
Diversity Level Definition
1: Limited Diversity The majority of prompts are near-identical or repeated with trivial modifications, such as basic rewordings (e.g., replacing “nude” with “nudity”). There is negligible lexical, structural, or conceptual diversity. Most prompts are variations on a fixed template and rely on the same narrow set of triggering keywords or phrases.
2$\sim$3: Moderate Diversity Prompts exhibit moderate diversity, often using modestly different triggering synonyms (e.g., replacing “nude” with “sexual”) or introducing light syntactic changes (e.g., changing the setting from “on the bed” to “in the room”). However, they still rely on a small group of core visual ideas and maintain similar structure and phrasing, with only minor surface-level differences.
4$\sim$5: High Diversity Prompts move beyond a small set of repetitive trigger words or formulaic expressions, demonstrating meaningful exploration of lexical, syntactic, and semantic alternatives. Instead of repeatedly relying on single terms like “nude woman,” the prompts vary across subjects (e.g., “erotic dancer,” “seductive character”), the frame (e.g., “softly lit room,” “posing suggestively”), and the scene structure. The prompts reflect creative and distributed discovery of diverse, or even unexpected potential triggers for generating unsafe content.
Appendix C More Discussion
Discussion with Other Optimization Methods. In addition to our proposed GC-SPSA algorithm, other optimization approaches can also be considered for minimizing the final loss function (i.e., Eq. (5)). One such alternative is reinforcement learning (RL) [51, 50], which are recently introduced to red team LLMs. A representative example is the recent CRT framework [50], which leverages PPO combined with a curiosity-driven reward and entropy bonus to enhance prompt diversity. Although CRT was originally designed for LLMs, the appendix of their revised paper preliminarily show that the same RL-based method can be extended to red team the Vanilla SD v1.5. However, CRT as implemented in the original paper only targets a general “unsafe” category. Therefore, it is not directly comparable to our method and other baselines, which is designed for specific unsafe categories. To enable a fairer comparison, we adapt CRT for a category-specific case study focusing on the “sexual” category. In this adapted version, the LLM is explicitly instructed to generate prompts related to sexual content. We train this CRT variant for 5,000 iterations, which takes approximately 24 hours on two NVIDIA RTX A100 GPUs, using all other default hyperparameters. As shown in Tab. XII, CRT exhibits high instability and, in some settings such as ESD, even fails completely. Moreover, in certain runs on CA, we observe CRT converges to repeatedly generating highly similar prompts after around 4,000 training steps, highlighting a lack of diversity and instability. In contrast, our method, DREAM, does not exhibit this repetitive behavior. We hypothesize that these differences arise from the fundamentally different optimization strategies used by RL and SPSA. In RL, each prompt (i.e., policy) is directly evaluated and rewarded. For safety-aligned models where most prompts fail to succeed, a few effective prompts receive disproportionately high rewards. This encourages the model to repeatedly generate the same high-reward prompts, even when regularized with KL divergence and novelty rewards, thereby reducing output diversity. In contrast, GC-SPSA optimizes by perturbing parameters and estimating the gradient via batch-based Monte Carlo sampling, without directly associating rewards with individual prompts. This means the model learns which direction in the parameter space improves performance rather than which specific prompts perform best. As a result, it is less likely to overfit to a single high-performing prompt, maintaining both stability and diversity throughout training. Nevertheless, we note that this analysis is not intended to diminish the potential of reinforcement learning. RL remains a powerful and expressive optimization paradigm, and with carefully tuned hyperparameters or alternative reward shaping strategies, its performance could potentially be improved. Recent advances such as GRPO [84], for instance, suggest promising directions that may help mitigate some of the issues that CRT suffer from. Our core claim is not that zeroth-order optimization is categorically superior, but rather that even a relatively simple and lightweight method like our GC-SPSA can already achieve stable and effective results in this challenging red teaming setting. We leave deeper exploration and more extensive optimization strategies to future work.
TABLE XII: Performance of CRT [50]. The results show the mean and standard deviation for 3 independent runs.
Metric SD v1.5 CA ESD SC
PSR $\uparrow$ 93% ± 06% 47% ± 14% 04% ± 03% 12% ± 08%
PS $\downarrow$ 0.65 ± 0.03 0.68 ± 0.04 0.72 ± 0.02 0.71 ± 0.01
Ethics Considerations. This research is intended solely for advancing the safety of T2I generative systems. DREAM is designed as a red teaming framework to proactively evaluate and improve existing safety mechanisms, not to attack or undermine any deployed systems. Nonetheless, we acknowledge that our method may generate prompts capable of bypassing safety mechanisms and elicit harmful outputs. To mitigate risk, we have responsibly disclosed all discovered vulnerabilities including prompt examples and generated images to the developers of affected models, safety filters, and commercial platforms via their official contact channels (e.g., forums and official emails), and are currently awaiting their responses. We have comprehensively discussed this work, including the user study, with our Institutional Review Board (IRB) and received an exempt determination. All study participants were adults in good physical and mental health, and confirmed to have no histories of heart conditions, psychological disorders, or vasovagal syncope, provided informed consent, and were made aware of their right to withdraw at any time. No personal or sensitive data was collected. Additionally, we place a strong emphasis on the well-being of both the researchers and study participants involved in this project. All authors and participants involved in this study were provided access to institutional mental health resources and encouraged to monitor and communicate any discomfort arising from interacting with unsafe or disturbing content during experimentation. Study participants were similarly given appropriate content warnings and support resources. We believe that fostering a research culture of care and accountability is essential, especially when dealing with high-stakes safety and content moderation work. To minimize potential misuse, we have refrained from publicly releasing the full set of discovered unsafe prompts until the affected parties have had sufficient time to address the issues. We will continue to engage with stakeholders and iterate on responsible disclosure practices in line with the principles outlined in the Menlo Report and the broader computer security research community.